# Modelling the interaction between ethnicity and infectious disease transmission dynamics

### v1

@@ 3.2.1 Specified $\epsilon$
We now investigate the inclusion of assortativity in the model to represent preferential inter-ethnicity interactions. There was no immediately obvious choice for the value of assortativity, so we first explored model outputs over the full range of $0\leq\epsilon\leq 1$ (plots of this are present in Appendix B.). We then estimated the value for assortativity and the transmission rates (Figure 4) using the statistical area data. We present the results for the SA2-prioritised fit here; results for SA1/SA2 total ethnicity are present in Appendix C. The transmission rates fit using the SA1-total and SA2-total data were qualitatively similar, but the SA1-total data had the highest values for $\epsilon$. This was an expected result due to smaller subdivisions sampling a smaller number of people, which increases the assortativity we would expect.
3.2.2 Non-parametric mixing
@@
We now construct the non-parametric transmission matrix directly from the population data, without imposing a particular mixing pattern to eliminate assumptions about national proportional mixing and assortativity. Again, the constructed matrix was fit such that the attack rates from the model match the data. The transmission rates can be seen in Figure 4. When compared to the proportionate mixing rates, the non-parametric rates for Māori and Pacific rates were lower, while the Asian and European/Other rates were higher, which meant there was a lesser disparity between the rates.
@@
Figure 4: The transmission rates fit to an a) assortative model and b) non-parametric model. The transmission rates were fit such that the attack rates from the model matched the attack rates from the data adjusted with the four CARs considered.
@@
We compared the transmission matrices for the assortative mixing model with the non-parametric model in Figure 5. The matrices remained relatively similar; the elements in the non-parametric matrix were 0.65-1.35 times the values in the assortative matrix, with the biggest difference between methods being the frequency of interaction between the Pacific and the Asian and European/Other groups, which had the biggest relative decrease and increase, respectively. All Pacific interactions increased except for the Pacific-European/Other interaction, which was substantially lower. The diagonals changed less than the off-diagonal elements, with the non-parametric diagonal elements being between 0.91 and 1.20 times the assortative matrix elements (with intra-Pacific interactions relatively increasing the most and intra Māori interactions relatively decreasing the most). The transmission rates and initial reproduction numbers remained similar.
@@
Figure 5: A comparison between a) the proportionate mixing transmission matrix; b) the assortative transmission matrix ($\epsilon=0.179$); c) the non-parametric transmission matrix; d) the proportionate mixing per capita transmission matrix; e) the assortative per capita transmission matrix; and f) the non-parametric per capita transmission matrix. Matrices were fit with the prioritised SA2 population data (where relevant) such that the attack rates from the model matched the attack rates from the case data with an assumed CAR of 50%.
@@
Table of transmission rate values for prioritised SA2. $a_{k}$ CAR $\epsilon$ Māori Pacific Asian European/Other Proportionate Mixing 4060 0.589 1.001 0.230 0.261 40 0.528 0.899 0.344 0.396 50 0.474 0.738 0.325 0.371 60 0.446 0.669 0.314 0.357 Assortative Mixing 4060 0.150 0.578 0.935 0.250 0.282 40 0.173 0.520 0.840 0.358 0.407 50 0.180 0.465 0.686 0.339 0.382 60 0.183 0.437 0.620 0.327 0.367 Non-parametric mixing 4060 0.584 0.859 0.250 0.300 40 0.520 0.802 0.353 0.416 50 0.465 0.658 0.333 0.388 60 0.436 0.597 0.321 0.373
@@ 3.2.3 Scenario analysis
Attack rates varied to differing degrees as a result of changes to model parameters. When setting $\epsilon$ to zero or setting each group’s vaccination rate to the population average, there was a relatively small change to the attack rates (Figure 6). Conversely, when setting the transmission rates to the population average, there was a significant change to the attack rates; this showed the ethnic disparity in the number of infections was dominated by the disparity in transmission rates and not the assortativity value nor the disparity in vaccination rates. It is worth differentiating the effect of vaccines and the effect of the disparity in vaccination. Vaccines did have an effect on reducing incidence rate and, in the real world, substantially reduced mortality and severity of infection. The difference in vaccination coverage over ethnic groups had a relatively smaller effect on differences in attack rates by ethnicity compared to transmission rate differences. Similar results hold for total measure statistical area data, present in Appendix C.
Figure 6: Comparison between the attack rates of the assortative mixing model with 1) no changes, 2) $\epsilon$ set to zero, 3) vaccination rates averaged, and 4) transmission rates averaged (see Table 2.3). The assortative matrix with no changes was fit to match the attack rates from the data with prioritised SA2 data and an assumed CAR of 50%. This meant the attack rates of 1) in the plot match the data. When the parameters were altered, the values were not refit.
@@ 4 Discussion
We explored the first Omicron wave during the COVID-19 pandemic in Aotearoa using an ethnicity-stratified SEIR model. We fit ethnicity-specific transmission rates and an assortativity constant such that the model’s attack rates matched those from the COVID-19 case data using three main assumptions of mixing: proportionate, assortative, and non-parametric mixing. We then attempted to quantify the relative importance of different drivers in the model.
Comparison of different scenarios revealed that, in our model, disparities in vaccination rates only explain a relatively small part of the observed differences in attack rates, so other factors must have been more important contributors. Likewise, assortative mixing alone is insufficient to explain these differences and can similarly only explain a relatively small difference. Differences between ethnicities in their average transmission rates explained a relatively large part of the observed differences, and therefore transmission rates are likely an important contributor. This is one possible explanation that is consistent with the observed attack rate disparity.
To reproduce the ethnicity-specific attack rate data from Aotearoa’s first Omicron wave with an SEIR model required substantial differences in transmission rates between ethnicities. These transmission rates represent differences in social interactions and not biological responses, i.e., there is not a biological difference between ethnicities. Māori, Pacific Peoples, and Asians had transmission rates 1.28, 1.99, and 0.88 times higher than New Zealand Europeans under the assumption of proportionate mixing. This result was consistent with findings from Ma et al.; however, the relative differences in our transmission rates were slightly lower than for Ma et al. (2.25, 1.62, 0.86, and 1.28 times the non-Hispanic white rate in New York City for Hispanics or Latinos, non-Hispanic Black people, non-Hispanic Asians, and multiracial or other people, respectively) [23]. Like Ma et al., we found that relative differences between contact rates were smaller under the assumption of assortative mixing as opposed to proportionate mixing. $\epsilon$ was significantly lower than the Ma et al. value, being 18.0% instead of 46% in New York City. Reasons for these differences in the Aotearoa context, covered in the next paragraphs, could include workplace, average household occupancy, or age-distribution effects.
Work by Chang et al. has shown higher infection rates among disadvantaged racial and socioeconomic groups in part due to their inability to self-isolate as quickly as other groups [36]. This was partly due to workplace effects. Māori and Pacific Peoples were more likely to work in sectors or roles such as agriculture, transport, and manufacturing jobs [37]. While there were no lockdowns during the Omicron wave, these businesses would be operating close to normal levels during the first Omicron wave and would have less opportunity to work from home compared to other sectors. Risk would have also been increased working jobs that are more person-facing. A combination of these factors could have increased the risk and number of interactions Māori and Pacific Peoples faced during the first wave of Omicron.
Māori have historically been disadvantaged socio-economically, leading to worse health outcomes [38]. This disparity could be contributing to the disparity in transmission rates. In Switzerland, in 2021, adults and young adults living in higher socio-economic positions had fewer contacts than those in lower socio-economic positions. [39]. This would suggest socio-economic status could partly explain observed differences between ethnicities. Māori and Pacific Peoples typically have higher levels of deprivation that lead to an increased risk of hospitalisations from diseases such as influenza [40]. Deprivation is not only associated with the severity of infection but also the incidence rate of diseases. People who live in deprived areas are more likely to live in households with more people in them. Household crowding is a known risk factor for infectious diseases, like COVID-19, and disproportionately affects Māori, Pacific Peoples, and Asians [41, 42]. In 2013, 20% of Māori, 38% of Pacific Peoples, and 18% of Asian people lived in crowded households, compared to just 4% of Europeans. In China, 75%-80% of clustered infections were within families, showing high intra-family transmission rates [34]. A meta-analysis of studies has estimated that the risk of household transmission is ten times that of other contacts [43]. Higher household crowding was one factor that probably contributed to higher transmission rates for the Māori and Pacific groups. The picture is less clear for the Asian group, as despite the increased rate of household crowding for Asians, our model estimated lower transmission rates. One possible reason for this was adherence to COVID-19 response measures in which Asians have been shown to have increased adherence to response measures such as contact tracing and mask wearing, both in Aotearoa [44, 45] and overseas [46].
Māori and Pacific Peoples have a younger age distribution than Asian people and European/Others [47]. As a result, the average age of a case identifying as Māori will be substantially lower than for European cases. Some studies have estimated that the majority of household transmissions begin with a child [48], and young adults are frequently estimated to have higher contact rates than older people [49, 50]. This could contribute towards Māori and Pacific Peoples having higher transmission rates than European people. "Susceptibility to COVID-19 in individuals under 20 years of age has been estimated as half that of those over 20, which would mean Māori are expected to have fewer infections than European/Others. Another possible issue is that younger cases get reported less, as younger people manifest clinical symptoms less [22]. This could have led to fewer cases reported in younger populations, which would disproportionately affect Māori and Pacific Peoples and imply a lower CAR in these groups than for European/Other people.
In addition, other possible but less documented reasons for differences in transmission rates include differences in frequency of community interaction, method of transportation, and the cultural importance of interacting with extended family, which may be context-specific or hard to quantify. One example of cultural practices that could affect transmission rates was tangihanga, a Māori funeral practice, which contains a customary extended ceremony. During the pandemic, these ceremonies had measures to restrict COVID-19 spread, and at some points during the pandemic, tangihanga were not able to be held in any capacity [51]. During the Omicron wave, no more than 100 people were allowed into a marae, a traditional Māori tribal meeting place, at any one time. Strict COVID-19 measures were used at some marae to prevent transmission, despite this, tangihanga were still potential transmission events during Omicron [52]. As cultural activities were altered during COVID-19, it can be difficult to estimate the effect of these activities on disease transmission.
@@ 4.1 Sensitivity analysis
To analyse the sensitivity of our model, we varied the assumptions and data sources we used. The choice of CAR, data source for spatially disaggregated population sizes by ethnicity, and contact matrix method did not change the qualitative results from fitting the model. The value of CAR determines the magnitude of disparity but does not change the fact that disparity is present. A lower assumed CAR rate does lead to a greater disparity in the transmission rate, and when Māori and Pacific Peoples had a lower assumed CAR, the disparity in transmission rates was largest. Changing the assumed CAR did not change the relative ordering of transmission rates, i.e., Pacific Peoples always had the highest transmission rate. The values of transmission rates were not highly sensitive to whether we used population data from the Te Whatu Ora dataset or StatsNZ. These results were also not highly sensitive to whether we used SA1 or SA2 data. This shows that although quantitative estimates of model parameters differ, the qualitative conclusions are robust to the different data sources and assumptions tested.
@@ 4.2 Limitations
Our model ignores age-related effects and their interaction with ethnic effects, an obvious avenue for future work, which (as discussed above) could explain some of the observed disparities in attack rates. The same is true of combining other factors with ethnicity, such as deprivation/socioeconomic status. This is part of a larger limitation that we cannot definitively disentangle differences in transmission rates from differences in reporting rates or other factors. The main difference that drives disparity in COVID-19 cases is likely to be transmission rates, but there could be other factors that are not included in the model. This is made more difficult to interpret by the fact that these ethnicity-specific transmission rates are proxies for a range of socioeconomic and cultural factors that affect contact patterns and transmission risk. We do not model a causation effect but a correlation between ethnicity and infection rate. These factors can change in complicated ways over time, which means estimates of contact patterns are restricted to the specific context of the first Omicron wave and may not generalise to other situations, e.g. non-pandemic periods or a future pandemic.

### v2

@@ 4.2 Assortative mixing model
We now investigate the inclusion of assortativity in the model to represent preferential inter-ethnicity interactions. There was no immediately obvious choice for the value of assortativity, so we first explored model outputs over the full range of $0\leq\epsilon\leq 1$ (plots of this are presented in Supplementary Figure 1). We then estimated the value for assortativity and the transmission rates using the statistical area data (Figure 3). We present the results for the SA2-prioritised data here; results for SA1/SA2 total ethnicity data are presented in the Supplementary Figure 5. The transmission rates estimated using the SA1-total and SA2-total data were qualitatively similar, but the SA1-total data had the highest value for $\epsilon$. This was an expected result due to smaller subdivisions sampling a smaller number of people, which increases the assortativity we would expect.
4.3 Unconstrained mixing model
@@
We now construct the unconstrained mixing transmission matrix directly from the population data, without imposing a particular mixing pattern to eliminate assumptions about national proportional mixing and assortativity. Again, the transmission rates were chosen such that the attack rates from the model exactly match the data. The transmission rates can be seen in Figure 3. When compared to the proportionate mixing rates, the unconstrained mixing model rates for Māori and Pacific rates were lower, while the Asian and European/Other rates were higher, which meant there was a lesser disparity between the ethnic transmission rates.
@@
Figure 3: a) Transmission rates fit to the proportionate mixing model, b) the reproductive number of the proportionate mixing model, c) transmission rates fit to the assortative mixing model and d) transmission rates fit to the unconstrained mixing model. The transmission rates were chosen such that the attack rates from the model exactly matched the attack rates from the data, after adjustment with the assumed CARs. SA2 prioritised data was used for fitting the assortative mixing and unconstrained mixing models.
@@
We compared the transmission matrices for the assortative and the unconstrained mixing models in Figure 4. The matrices were relatively similar; the elements in the unconstrained mixing model matrix were 0.65-1.38 times the values in the assortative matrix, with the biggest difference between models being the frequency of interaction between the Pacific and the Asian and European/Other groups. All Pacific interactions were higher under the unconstrained model than the assortative model except for the Pacific-European/Other interaction, which was substantially lower. The diagonal elements differed less between the two models than the off-diagonal elements, with the unconstrained model diagonal elements being between 0.87 and 1.19 times the assortative matrix elements (with intra-Pacific interactions relatively increasing the most and intra-Māori interactions relatively decreasing the most). The transmission rates and initial reproduction numbers were similar for both models.
Figure 4: A comparison between a) the proportionate mixing model transmission matrix; b) the assortative mixing model transmission matrix ($\epsilon=0.179$); c) the unconstrained mixing model transmission matrix; d) the proportionate mixing model per capita transmission matrix; e) the assortative mixing model per capita transmission matrix; and f) the unconstrained mixing model per capita transmission matrix; g) the difference between the proportionate and assortative mixing model per capita matrix; h) the difference between the proportionate and unconstrained mixing model per capita matrix; i) the difference between the assortative and unconstrained mixing model per capita matrix. Per capita matrix values are all scaled by $10^{-7}$
@@
. Matrices were fit with the prioritised SA2 population data (where relevant) such that the attack rates from the model exactly matched the attack rates from the case data with an assumed CAR of 50%.
@@
Table of transmission rate values for prioritised SA2. $a_{k}$ CAR $\epsilon$ Māori Pacific Asian European/Other Proportionate Mixing 40/60 0.633 1.098 0.257 0.302 40 0.569 0.988 0.386 0.461 50 0.510 0.805 0.364 0.430 60 0.478 0.727 0.351 0.412 Assortative Mixing 40/60 0.145 0.619 1.027 0.279 0.325 40 0.176 0.557 0.920 0.402 0.475 50 0.183 0.497 0.746 0.379 0.443 60 0.186 0.466 0.672 0.365 0.424 Unconstrained mixing 40/60 0.626 0.938 0.279 0.347 40 0.558 0.877 0.396 0.485 50 0.498 0.714 0.372 0.451 60 0.466 0.645 0.359 0.431
4.4 Sensitivity analysis
To analyse the sensitivity of our model, we varied the assumptions and data sources we used (present in Supplementary Table 2). The choice of CAR, data source for spatially disaggregated population sizes by ethnicity, and mixing model did not change the qualitative results from fitting the model. The value of CAR determines the magnitude of ethnic disparity but does not change the fact that disparity is present. When we assumed a lower value for CAR rate, it lead to a greater disparity in the ethnicity-specific transmission rates, and when Māori and Pacific Peoples had a lower assumed CAR, the disparity in transmission rates was largest. Changing the assumed CAR did not change the relative ordering of transmission rates, i.e., Pacific Peoples always had the highest transmission rate. These results were also not highly sensitive to whether we used SA1 or SA2 data.
@@ 4.4.1 Scenario analysis
Attack rates varied to differing degrees as a result of changes to model parameters representing the scenarios defined in Table 3.3. When setting $\epsilon$ to zero (scenario 2) or setting each group’s vaccination rate to the population average (scenario 3), there was a relatively small change to the model attack rates (Figure 5). Conversely, when setting the transmission rates to the population average (scenario 4), there was a significant change to the attack rates; this showed the ethnic disparity in the number of infections was dominated in the model by the disparity in ethnicity-specific transmission rates and not the assortativity value nor the disparity in ethnic vaccination rates. Note that we have only investigated the effect of disparities in vaccination rates on infections. It is likely that disparities in vaccination rates were responsible for larger inequities on mortality and morbidity due to the higher vaccine effectiveness against sever disease and death than against infection or transmission.
Figure 5: Comparison between the attack rates of the assortative mixing model with 1) ethnicity-specific parameters and fitted value of $\epsilon$, 2) scenario 1 with $\epsilon$ set to zero, 3) scenario 1 with population-averaged vaccination rates, and 4) scenario 1 with population averaged transmission rates (see Table 3.3). The assortative model ethnicity-specific parameters and $\epsilon$ were fit to match the attack rates from the prioritised SA2 data with an assumed CAR of 50%. This meant the attack rates of scenario 1 matched the attack rates found from the data.
@@ 5 Discussion
We explored the first Omicron wave during the COVID-19 pandemic in Aotearoa using an ethnicity-stratified SEIR model. We estimated ethnicity-specific transmission rates and an assortativity constant such that the model’s attack rates matched those from the COVID-19 case data using three different mixing models: proportionate mixing, assortative mixing, and unconstrained mixing. We then attempted to quantify the relative importance of different drivers in the model.
@@ 5.1 Sensitivity Analysis
Our model was not sensitive to the spatial resolution of the population data (choice of SA1 or SA2 area level) used to estimate the parameters of the assortative mixing and unconstrained mixing models. Qualitative results did not change when varying the choice of mixing model or assumed value of CAR. This showed that although quantitative estimates of model parameters differ, the qualitative conclusions are robust to the different data sources and assumptions tested.
5.2 Mixing model comparison
All three mixing models fit the cumulative attack rate data exactly because we were using four data points (attack rates in each ethnicity group) to estimate four unknown parameters (transmission rates in each ethnicity group). The assortative mixing model had one additional unknown parameter (the assortativity constant), which was informed by the area-level population data. The mixing models presented make different assumptions about between-group and within-group contact rates. We consider the unconstrained mixing model to be the most representative of real-world contacts, as it is informed directly by empirical data and does not impose any assumptions on the structure of contact matrix. Assortative mixing gave a good approximation for this matrix, however the assumption of a population wide assortativity value was limiting. Proportionate mixing is significantly less representative of real-world mixing. This can be seen in ethnic distributions over Aotearoa being heavily clustered for Pacific and Asian groups (presented in Supplementary Figures 2 and 3).
5.3 Scenario analysis
Comparison of different scenarios revealed that, in our model, disparities in vaccination rates between ethnic groups only explain a relatively small part of the observed differences in the ethnicity-specific attack rates, so other factors must have been more important contributors. Likewise, assortative mixing alone is insufficient to explain these differences and can similarly only explain a relatively small difference. Differences between ethnicities in their average transmission rates explained a relatively large part of the observed differences in their attack rates, and therefore transmission rates are likely an important contributor. This is one possible explanation that is consistent with the observed ethnic attack rate disparity.
5.4 Transmission rate differences
To reproduce the ethnicity-specific attack rate data from Aotearoa’s first Omicron wave with an SEIR model required substantial differences in transmission rates between ethnicities. These transmission rates represent differences in social interactions and not biological responses, i.e., there is not a biological difference between ethnicities. The Māori, Pacific, and Asian transmission rates were estimated to be 1.19, 1.87, and 0.85 times higher than European/Other rate under the assumption of proportionate mixing. The approximate magnitude of the relative differences in transmission rates between groups was consistent with findings from Ma et al. [ma_modeling_2021] for New York in 2020; however, the relative differences in our transmission rates were slightly lower than for Ma et al. (2.25, 1.62, 0.86, and 1.28 times the non-Hispanic White rate in New York City for the Hispanic or Latino, non-Hispanic Black, non-Hispanic Asian, and multiracial or other groups, respectively) [ma_modeling_2021]. Like Ma et al., we found that relative differences between transmission rates were smaller under the assumption of assortative mixing as opposed to proportionate mixing. The estimated value of the assortativity constant $\epsilon$ was significantly lower than the Ma et al. value, being 18.3% instead of 46% in New York City. Reasons for these differences in the Aotearoa context, covered in the next section, could include workplace, average household occupancy, or age-distribution effects.
5.5 Possible reasons for differences
5.5.1 Jobs and self-isolation
Work by Chang et al. has shown higher infection rates among disadvantaged racial and socioeconomic groups in part due to their inability to self-isolate as quickly as other groups [chang_mobility_2021]. This was partly due to workplace effects. Māori and Pacific Peoples were more likely to work in sectors or roles such as agriculture, transport, and manufacturing jobs [harvey_summary_2023]. While there were no lockdowns during the Omicron wave, these businesses would be operating close to normal levels during the first Omicron wave and would have less opportunity to work from home compared to other sectors. Risk would have also been increased working jobs that are more person-facing. A combination of these factors could have increased the risk and number of interactions Māori and Pacific Peoples faced during the first wave of Omicron.
5.5.2 Deprivation and household crowding
Māori have historically been disadvantaged socio-economically, leading to worse health outcomes [robson_social_2007]. This disparity could be contributing to the disparity in ethnicity-specific transmission rates. In Switzerland, in 2021, adults and young adults living in higher socio-economic positions had fewer contacts than those in lower socio-economic positions. [domenico_individual-based_2025]. This would suggest socio-economic status could partly explain observed differences between ethnicities. Māori and Pacific Peoples typically have higher levels of deprivation that lead to an increased risk of hospitalisations from diseases such as influenza [lopez_influenza_2016]. Deprivation is not only associated with the severity of infection but also the incidence rate of diseases. People who live in deprived areas are more likely to live in households with more people in them. Household crowding is a known risk factor for infectious diseases, like COVID-19, and disproportionately affects Māori, Pacific, and Asian Peoples [ministry_of_health_analysis_2014, johnson_stocktake_2018]. In 2013, 20% of Māori, 38% of Pacific Peoples, and 18% of Asian Peoples lived in crowded households, compared to just 4% of Europeans. In China, 75%-80% of clustered infections were within families, showing high intra-family transmission rates [chen_fangcang_2020]. A meta-analysis of studies has estimated that the risk of household transmission is ten times that of other contacts [lei_household_2020]. Higher household crowding was one factor that probably contributed to higher transmission rates for the Māori and Pacific groups. The picture is less clear for the Asian group, as despite the increased rate of household crowding for Asian Peoples, our model estimated lower transmission rates. One possible reason for this was adherence to COVID-19 response measures in which Asian Peoples have been shown to have increased adherence to response measures such as contact tracing and mask wearing, both in Aotearoa [chambers_evaluation_2024, gray_wearing_2020] and overseas [hearne_understanding_2022].
5.5.3 Age distribution effects
Māori and Pacific Peoples have a younger age distribution than Asian people and European/Others [bryant_ageing_2003]. As a result, the average age of a case identifying as Māori will be substantially lower than for European cases. Some studies have estimated that the majority of household transmissions begin with a child [tseng_smart_2023], and young adults are frequently estimated to have higher contact rates than older people [prem_projecting_2017, hoang_systematic_2019]. This could contribute towards Māori and Pacific Peoples having higher transmission rates than European people. Susceptibility to COVID-19 in individuals under 20 years of age has been estimated as half that of those over 20, which would mean Māori are expected to have fewer infections than European/Others. Another possible issue is that younger cases get reported less, as younger people manifest clinical symptoms less [davies_age-dependent_2020]. This could have led to fewer cases reported in younger populations, which would disproportionately affect Māori and Pacific Peoples and imply a lower CAR in these groups than for European/Other people.
5.5.4 Other possible effects
In addition, other possible but less documented reasons for differences in transmission rates include differences in frequency of community interaction, method of transportation, and the cultural importance of interacting with extended family, which may be context-specific or hard to quantify. One example of cultural practices that could affect transmission rates was tangihanga, a Māori funeral practice, which contains a customary extended ceremony. During the pandemic, these ceremonies had measures to restrict COVID-19 spread, and at some points during the pandemic, tangihanga were not able to be held in any capacity [rangiwai_impacts_2021]. During the Omicron wave, no more than 100 people were allowed into a marae, a traditional Māori tribal meeting place, at any one time. Strict COVID-19 measures were used at some marae to prevent transmission, despite this, tangihanga were still potential transmission events during Omicron [cooper_covid_2022]. As cultural activities were altered during COVID-19, it can be difficult to estimate the effect of these activities on disease transmission.
@@ 5.6.1 Interaction with other factors
Our model ignores age-related effects and their interaction with ethnic effects, an obvious avenue for future work, which (as discussed above) could explain some of the observed ethnic disparities in attack rates. The same is true of combining other factors with ethnicity, such as deprivation/socioeconomic status. This is part of a larger limitation that we cannot definitively disentangle differences in transmission rates from differences in reporting rates or other factors. The main difference that drives ethnic disparity in COVID-19 cases is likely to be transmission rates, but there could be other factors that are not included in the model. This is made more difficult to interpret by the fact that these ethnicity-specific transmission rates are proxies for a range of socioeconomic and cultural factors that affect contact patterns and transmission risk. We do not model a causation effect but a correlation between ethnicity and infection rate. These factors can change in complicated ways over time, which means estimates of contact patterns are restricted to the specific context of the first Omicron wave and may not generalise to other situations, e.g. non-pandemic periods or a future pandemic.
5.6.2 Model fitting and parameters
In our model, we treated the first Omicron wave in Aotearoa as a SEIR epidemic that ignores differences in time-dependent case rates, which are not well explained by a SEIR model. Our model aimed to be as simple as possible and in this pursuit we used constant parameter values for the ethnicity-specific contact rates. Because our model only had four unknown parameters (transmission rates for the four ethnicity groups), we were able to choose these to provide an exact match with cumulative ethnicity-specific attack rate data. As a consequence, a single model fit does not provide any information about uncertainty; instead we approached this via a sensitivity analysis to investigate how model results were affected by alternative assumptions about mixing patterns, case ascertainment rates, and different choices of area-level population data.
We note that an alternative approach would be to incorporate additional flexibility in the model, via time-varying parameters or the inclusion of other model mechanisms. This would result in a higher-dimensional parameter inference problem, which could be tackled by a Bayesian computational inference method, such as Markov chain Monte Carlo or approximate Bayesian computation, using the full time series of case data as opposed to just the cumulative attack rates. This approach would have the advantage of incorporating parameter uncertainty and enabling systematic comparison of model fit across competing models. However, the drawback of this would be that it would require a much more complex model to obtain a reasonable fit with the time-dependent data, with multiple assumptions about the relevant mechanisms [datta_modelling_2025]. This could risk obscuring the dominant processes affecting observed trends, and suffer from parameter non-identifiability issues that are common in these types of model [roosa_assessing_2019]. Our simplified modelling approach needs to be interpreted with caution due to its limitations. Nevertheless, we believe that it has value as a broad-brush analysis that enables the approximate magnitude of disparities between ethnicity groups and the relative contributions of different effects to be compared in a simplified framework.